Question 6

Classify each of the following functions as algebraic or transcendental.

a) 𝑓(π‘₯) = tan(8π‘₯)

b) 𝑓(π‘₯) = √π‘₯5+3 9π‘₯+5

c) 𝑓(π‘₯) = 9π‘₯

d) 𝑓(π‘₯) = 1+π‘₯

A company that operates for 18 hours a day has introduced a new daily wage system that incorporates variable hourly rate to encourage longer shifts. The company pays $10 for first hour or any part of the first hour. Afterwards, company pays additional 2% per hour for each hour or part of each additional hour. Write a piece-wise function that shows the salary as a function of hours. Also sketch a graph of this function.

Question 8

Describe how the function 𝑓(π‘₯) = βˆ’(π‘₯ + 1)2 βˆ’ 4 can be graphed using the graph of 𝑦 = π‘₯2 and a sequence of transformations.


1
Expert's answer
2021-09-24T16:21:02-0400

"1. \text{ A transcendental function is an analytic function that does not satisfy a polynomial}\\
\text{equation while In mathematics, an algebraic function is a function that can be defined}\\
 \text{ as the root of a polynomial equation. The function a is}\text{ transcedental while the functions}\\\text{ b, c, d are algebraic }\\
2. \text{The piecewise function can be given by }\\
f(x) = 10 \text{ if $0 < x \leq 1 $}\\
f(x) = \frac{2}{100}.10(x-1)+10 \text{ if 1$\leq x \leq 18$}\\
\implies f(x) = \begin{cases} 10 \text{ if 0 < x $\leq$ 1 }\\
\frac{2x+98}{10} \text{ if $1< x \leq 18$}
 \end{cases}\\
3. \text{ First the function $y = x^2$ undergoes reflection by multiplying it by -1, it is then }\\
\text{shifted 1 unit to the right, and it is shifted 4 units downward to get}\\
y = -(x+1)^2 -4\\
\text{The transformations can be represented graphically in the following images}\\"